OpenStudy (anonymous):

the perimeter of a rectangle is 54ft. Twice the length is 3ft. more than the width. What is the size of the rectangle? Legnth_____ft. Width_____ft

OpenStudy (lgbasallote):

let the length be L twice the length is 3ft more than the width so... 2L = W + 3 <=== W is width the formula for perimeter is L + L + W + W = P if we substitute 2L = W + 3 here... L + L + (2L - 3) + (2L - 3) = P so if we substitute P = 54.... L + L + 2L - 3 + 2L - 3 = 54 combine like terms... 6L - 6 = 54 add 6 to both sides... 6L = 60 divide both sides by 6 L = 10 it says that 2L = W + 3 2(10) = W + 3 20 = W + 3 so W = 17 does that help?

OpenStudy (anonymous):

Thank You :)